OpenStudy (anonymous):

a twenty minute audio tape has four songs with times in minutes and seconds as follows:a, 4:30/ B, 4:00/ C, 5:30/ D, 6:00. If the tape is started at any random position, what is the probability that it will be playing song A or C?

OpenStudy (m):

50%

OpenStudy (anonymous):

are yoj making that up?

OpenStudy (m):

(4.5 minutes + 5.5 minutes) / (20 minutes) = 10/20 = 0.5 = 50%